OpenStudy (anonymous):

what is the ratio of the circumfrences for the two circles with areas 6pie m^2 and 150pie m^2?

OpenStudy (anonymous):

one to five. genebrowder will explain

OpenStudy (anonymous):

the ratio of the Circumference C= 1/5

OpenStudy (anonymous):

oh i guess you still need an explanation. the ratio of the areas is \[\frac{6\pi m^2}{152\pi m^2}=\frac{1}{25}\]

OpenStudy (anonymous):

the area is two dimensional, comes with the square. the circumference is one dimensional, so take the square root of that ratio. the square root of \[\frac{1}{25}\] is \[\frac{1}{5}\]

OpenStudy (anonymous):

A1=6pi A2=150pi A1/A2= 6pi/150pi=1/25

OpenStudy (anonymous):

for diameter D1/D2=sqrt(6pix4)/(sqrt150pix4)=2sqrt6/10sqrt6=1/5

OpenStudy (anonymous):

and also for circumference C1/C2=piD1/piD2=D1/D2= 1/5 or 1 is to 5

OpenStudy (anonymous):

is it a little bit understandable Made?

OpenStudy (anonymous):

i understood it a little while ago , but thank you very much for explaining it wayy out atleast now i can make sure